OpenStudy (anonymous):

Logical quantifiers x=5 => x is odd f(x) is continuous <= f(x) is differentiable The graph of f(x) never touches x-axis <=> f(x)=0 has no real solutoins x=1=>x^4=1 sin^2(x)=1/4<=> cos^2(x)=3/4

OpenStudy (anonymous):

One of these is wrong for some reason... i dont know which one is wrong

OpenStudy (anonymous):

x usually denotes a real number. So I would say if you're talking about REAL 4th roots then the second to last one would be iff?

OpenStudy (anonymous):

Or <=> because there is only 1 REAL 4th root to 1, if you don't consider complex. I just feel like complex are always denoted by z. So if the context is real numbers then its <=> imo.

OpenStudy (anonymous):

can x =-1?

OpenStudy (anonymous):

what i thought is that x^4=1 can => x=1 or x=-1 , but x=1 only yields x^4=1?

OpenStudy (anonymous):

Well it depends on how specific you want to be. TECHNICALLY speaking the definition of a square root is: Suppose s is a positive real number, then r is called the square root of s if r is a positive real number and r^2=s.

OpenStudy (anonymous):

So technically speaking for square roots the negative ones "don't count" but realistically speaking, negative numbers square to positive ones.

OpenStudy (anonymous):

yea i see where u are going

OpenStudy (anonymous):

Every non-negative real number x has a unique non-negative square root, called the principal square root, which is denoted by , where is called radical sign. For example, the principal square root of 9 is 3, denoted , because 32 = 3 × 3 = 9 and 3 is non-negative.

OpenStudy (anonymous):

Thats from wikipedia. Ignore where symbols are missing, just copy pasted.

OpenStudy (anonymous):

But its just how "technical" you wanna be?

OpenStudy (anonymous):

i dont know cuz this is one of my homework..so there were no clearfications..

OpenStudy (anonymous):

:/ I'm sorry, I hate things like that.